1. For a given binomial distribution with n fixed trials and p, which is the probability of success of each trial, the binomial distribution is skewed left if p=0.50

Select one:

True or False

2. Consider a graph of a normal distribution with the mean μμ and standard deviation σσ. The graph will never cross the horizontal axis. This happens because the normal distribution is an exponential function.

Select one:

True or False

3. Suppose you flip a fair a coin 20 times. You flip the coin the same way for each flip. The fair coin can land as a head or a tail. In other words, there are only two outcomes for each coin flip. You wish to find the probability that you will get twelve heads out of the 20 flips. This is an example of a binomial experiment

Select one:

True or False

4. Let P(A)=0.5P(A)=0.5 and P(B)=0.3P(B)=0.3. If P(A|B)=0.2P(A|B)=0.2, calculate P(AandB)P(AandB).

Select one:

a. 0.80

b. 0.60

c. 0.10

d. 0.06

5. Given P(A)=0.6P(A)=0.6 and P(B)=0.5P(B)=0.5, and P(A|B)=0.3P(A|B)=0.3. Calculate P(AorB)P(AorB).

Select one:

a. 0.80

b. 0.40

c. 0.50

d. 0.95
Answer #1

1. Answer: False

If p = 0.50 then binomial distribution is symmetric.

--

2. Answer: False

--

3. Answer : True

--

4. P(A)=0.5, P(B)=0.3, P(A|B)=0.2

P(A and B) = P(A|B)*P(B) = 0.2*0.3 = 0.06

Answer: d. 0.06

--

5. P(A)=0.6, P(B)=0.5, P(A|B)=0.3

P(A and B) = P(A|B)*P(B) = 0.3*0.5 = 0.15

P(A or B) = P(A) +P(B) -P(A and B) = 0.6 + 0.5 - 0.15 = 0.95

Answer: d. 0.95